Main bibliographic database about transient and IR low background photo detectors

A. Coulais

April 30, 2002

References

AAAC98 [1] A. Abergel and A. Coulais, New method to correct ISOCAM-LW transient effects : first results, tech. rep., IAS Lab., Septembre 1998.

ADA96 [2] A. Abergel, F.-X. Désert, and H. Aussel, IAS model for ISOCAM LW transient correction, technical report, Décembre 1996.

Abergel96 [3] A. Abergel et al., ISOCAM mapping of the ρ-Ophiuchi main cloud., Astron. Astrophys., 315:L329–L332, Novembre 1996.

AMDDPAS99 [4] A. Abergel, M. A. Miville-Deschêne, F.-X. Désert, M. Pérault, H. Aussel, and M. Sauvage, IAS model for ISOCAM LW transient correction, accepted in Experimental Astronomy, Février 1999.

AbergelPer94 [5] A. Abergel and M. Pérault, Flight model LW channel : noise due to transients, tech. rep., IAS lab., Juin 1994.

Acosta98 [6] J. A. Acosta-Pulido, C. Gabriel, and H. O. Castañeda, Transient effects in ISOPHOT data: status of medelling and correction procedures, tech. rep., ESA, 1998.

Agnese91 [7] P. Agnèse, J. J. Engelmann, and P. Mottier, Results of radiation tests performed on the ISOCAM infrared detector array, IEEE Trans. on Nuclear Science, 38:953–963, Août 1991.

Agnese89 [8] P. Agnèse, C. Lucas, P. Maillart, P. L. Mottier, Y. Lepennec, and P. Masse, The ISOCAM camera long-wavelength channel detector, Proceedings of SPIE, 1070:124–128, Juillet 1989.

Balleux2000 [9] F. Balleux, Réalisation d’une chaîne de traitement pour les données ISOPHOT C-100, tech. rep., Stage de Magistère, 2000.

Biviano98Dark [10] A. Biviano et al., The ISOCAM dark current calibration report, tech. rep., ISOCAM, Novembre 1998.

Biviano98Flat [11] A. Biviano et al., The ISOCAM flat field calibration report, tech. rep., ISOCAM, Novembre 1998.

BloGorKho89 [12] I. K. Blokhin, T. A. Gorshkova, and V. A. Kholodnov, Nonlinear distortions of the photocurrent of high-resistance impurity photoresistors, Scripta Technica, pp. 106–113, 1989.

BloKho86 [13] I. K. Blokhin and V. A. Kholodnov, Characteristics of the signal and noise currents in longitudinal extrinsic photoconductors in the background-limited regime, Sov. Phys. Semicond., 20(10):1208– 1210, 1986.

BOT92 [14] I. K. Blokhin, V. V. Osipov, and I. I. Taubkin, Coordinate dependence of photosensitivity and other pecularities of extrinsic photoconductor detectors at low backgrounds, Infrared Physics, 33(1):9– 26, 1992.

BloRaKho89 [15] I. K. Blokhin, A. A. Rakhubovskii, and V. A. Kholodnov, Frequency characteristics of an isotypic contact with a high-resistivity semiconductor, Sov. Phys. Semicond., 23(10):1082–1085, 1989.

Blommaert98 [16] J. Blommaert, ISOCAM photometry report, tech. rep., ESA Vilspa, Spain, Mai 1998.

Bouaissi99 [17] S. Bouaissi, étude des différents modèles de phénomènes transitoires sur le satellite ISO, tech. rep., Stage de license, 1999.

Bourot98 [18] N. Bourot, Imagerie en astronomie : correction des transitoires de la caméra ISOCAM, tech. rep., Stage de magistère, 1998.

Caux99 [19] E. Caux, Application of Fouks-Schubert transient correction method to LWS data: status and plans, tech. rep., CESR, CNRS, France, 1999.

Cesarsky92 [20] C. J. Cesarsky, ISOCAM, in Infrared Astronomy with ISO. L’astronomie infrarouge et la mission ISO., T. Encrenaz and M. Kessler, eds., Les Houches, France, June 13-27, 1991, Nova Science, Commack, N.Y., 1992, pp. 31–50.

Cesarsky94 [21] C. J. Cesarsky et al., Camera of the Infrared Space Observatory, Optical Engineering, 33:751–761, Mars 1994.

Cesarsky96 [22] C. J. Cesarsky et al., ISOCAM in flight., Astron. Astrophys., 315:L32–L37, Novembre 1996.

Church92ESAa [23] S. E. Church, M. J. Griffin, P. A. R. Ade, M. C. Price, R. J. Emery, and B. M. Swinyard, Performance testing of doped-germanium photoconductors for the ISO Long-Wavelength Spectrometer, in ESA SP-356 Photon Detectors for Space Instrumentation, Décembre 1992, pp. 255–260.

Church93 [24] S. E. Church, M. J. Griffin, M. C. Price, P. A. Ade, R. J. Emery, and B. M. Swinyard, Performance characteristics of doped-Ge photoconductors for the Infrared Space Observatory Long Wavelength Spectrometer, Proceedings of SPIE, 1946:116–125, Octobre 1993.

Church92ESAb [25] S. E. Church, M. C. Price, M. J. Griffin, P. A. R. Ade, R. J. Emery, and B. M. Swinyard, Non-linear effects in doped-germanium photoconductors for the ISO Long Wavelength Spectrometer, in ESA SP-356 Photon Detectors for Space Instrumentation, Décembre 1992, pp. 261–264.

Church96 [26] S. E. Church, M. C. Price, N. M. Haegel, M. J. Griffin, and P. A. R. Ade, Transient response in doped germanium photoconductors under very low background operation, Applied Optics, 35:1597–1604, 1996.

Clegg96 [27] P. E. Clegg et al., The ISO Long-Wavelength Spectrometer., Astron. Astrophys. Suppl. Ser., 315:L38–L42, Novembre 1996.

Coulais98modele [28] A. Coulais, Étude des équations de Fouks, tech. rep., IAS lab., Octobre 1998.

Coulais99sources [29] A. Coulais, État de l’étude des transitoires LW-ISOCAM sous fort gradient d’illumination, tech. rep., IAS lab., Mai 1999.

ACAA98 [30] A. Coulais and A. Abergel, Transient correction of the ISOCAM data with the Fouks-Schubert model: first results, tech. rep., IAS Lab., Septembre 1998.

ACAA98Conf [31] A. Coulais and A. Abergel, Transient correction of the LW-ISOCAM data with the Fouks-Schubert model : first results, in ISO Conference, Paris, ESA, ESA, Octobre 1998.

ACAA99Gretsi [32] A. Coulais and A. Abergel, Correction des transitoires de la caméra infrarouge du satellite ISO, vol. T-1, GRETSI’99, Septembre 1999, pp. 147–150.

ACAA2000 [33] A. Coulais and A. Abergel, Transient correction of the LW-ISOCAM data for low contrasted illumination, Astron. Astrophys. Suppl. Ser., 141:533–544, Février 2000.

CoulaisFouks2000 [34] A. Coulais, B. I. Fouks, J.-F. Giovannelli, A. Abergel, and J. Sée, Transient response of IR detectors used in space astronomy : what we have learned from ISO satellite, in Infrared Spaceborne Remote Sensing VIII, M. Strojnik and B. Andresen, eds., vol. 4131, San Diego, USA, Août 2000, SPIE, pp. 205–217.

DeGraauw2000 [35] T. de Graauw, Space Infrared/Submm Telescopes: Past and Future, Advances in Space Research, 25:2145–2157, 2000.

DeGraauw96 [36] T. de Graauw et al., Observing with the ISO Short-Wavelength Spectrometer., Astron. Astrophys. Suppl. Ser., 315:L49–L54, Novembre 1996.

Delattre96 [37] C. J. Delattre, About transients, the fitting method, technical report, Juillet 1996.

Delettrez94 [38] C. Delettrez, M. J. Griffin, C. Le Naour, S. Church, and J. Chabaud, The detector signal processing chain of the ISO Long-Wavelength Spectrometer, Experimental Astronomy, 4:213–235, 1994.

Desert99tech [39] F.-X. Désert, A modified IAS model for ISOCAM LW transient correction, tech. rep., LAOG, Mars 1999. Version 2.0.

Desert99 [40] F.-X. Désert et al., A classical approach to faint extragalactic source extraction from ISOCAM deep surveys. Application to the Hubble Deep Field, Astron. Astrophys. Suppl. Ser., 342:363–377, Février 1999.

CAMhand99 [41] ESA, ISO Handbook Volume III (CAM), Version 1.0, Juillet 1999.

Fazio98 [42] G. G. Fazio, J. L. Hora, S. P. Willner, J. R. Stauffer, M. L. Ashby, Z. Wang, E. V. Tollestrup, J. L. Pipher, W. J. Forrest, C. R. McCreight, S. H. Moseley, W. F. Hoffmann, P. Eisenhardt, and E. L. Wright, Infrared array camera (IRAC) for the Space Infrared Telescope Facility (SIRTF), Proceedings of SPIE, 3354:1024–1031, Août 1998.

1997SPIE.3155...78F [43] A. M. Filachev, B. I. Fouks, and D. Greenfeld, Nature of distortions of the surface pattern produced by precise electron-beam processing of insulators, Proceedings of SPIE, 3155:78–88, Septembre 1997.

Fouks81a [44] B. I. Fouks, Injection properties of contacts to high-resistivity semiconductors. I, Sov. Phys. Semicond., 15(9):974–980, Septembre 1981.

Fouks81 [45] B. I. Fouks, Injection properties of contacts to high-resistivity semiconductors. I–II, Sov. Phys. Semicond., 15(9):974–980, 980–986, Septembre 1981.

Fouks81b [46] B. I. Fouks, Injection properties of contacts to high-resistivity semiconductors. II, Sov. Phys. Semicond., 15(9):980–986, Septembre 1981.

Fouks92ESA [47] B. I. Fouks, Nonstationary behaviour of low background photon detectors, in ESA SP-356 Photon Detectors for Space Instrumentation, Décembre 1992, pp. 167–174.

1993SPIE.1762..519F [48] B. I. Fouks, Nonstationary processes in extrinsic photoconductors, Proceedings of SPIE, 1762:519– 530, Janvier 1993.

Fouks94 [49] B. I. Fouks, On improvement of ISOPHOT detector operation, results of the visit at MPIA in 1994, tech. rep., MPIA, Mars 1994.

Fouks95theo [50] B. I. Fouks, On problems of operation of low-background IR detectors, Proceedings of SPIE, 2553:489–500, Septembre 1995.

Fouks1996modes [51] B. I. Fouks, Active modes of operation of low-background extrinsic IR detectors, Proceedings of SPIE, 2790:38–46, Avril 1996.

Fouks96 [52] B. I. Fouks, Phenomena in low-background IR detectors under high-energy particles, Proceedings of SPIE, 2817:160–171, Octobre 1996.

Fouks97 [53] B. I. Fouks, Theory of photoresponse of low-background IR detectors, Proceedings of SPIE, 3122:441– 452, Octobre 1997.

Fouks2001 [54] B. I. Fouks, Transient processes in high-field low-background IR detectors under inhomogeneous irradiation. I General statement of the problem, II Application to the ISOCAM detector array, tech. rep., IAS, Décembre 2001.

1994SPIE.2269..727F [55] B. I. Fouks, V. F. Kocherov, I. I. Taubkin, L. A. Vinokurov, and N. B. Zaletaev, Photoelectric cross talk between the elements of an extrinsic photoconductor detector array at low backgrounds, Proceedings of SPIE, 2269:727–735, Octobre 1994.

Fouks94Opt [56] B. I. Fouks, V. F. Kocherov, I. I. Taubkin, L. A. Vinokurov, and N. B. Zaletaev, Photoresponse of low-background extrinsic infrared detector arrays, Optical Engineering, 33:1485–1491, Mai 1994.

FouksSchu95 [57] B. I. Fouks and J. Schubert, Precise theoretical description of photoresponse for detectors of ISOPHOT’s Si:Ga array, Proceedings of SPIE, 2475:487–498, Juin 1995.

Fuji95 [58] M. Fujiwara, N. Hiromoto, and K. Araki, Moderate-fast transient response of the Ge:Ga photoconductors, Proceedings of SPIE, 2552:421–429, Septembre 1995.

Gabriel99 [59] C. Gabriel and J. A. Acosta-Pulido, The ISOPHOT interactive analysis (pia), The Universe as Seen by ISO. Eds. P. Cox & M. F. Kessler. ESA-SP 427., 427:73–76, Mars 1999.

IPAC3 [60] K. Ganga, G. Helou, R. Hurt, M.-S. Kong, J. Li, E. Valjavec, and D. Van Buren, ISO/CAM transient removal at IPAC, tech. rep., IPAC Lab., Août 1998.

IPAC2 [61] K. Ganga, M. Kong, L. Jing, and D. Van Buren, Transient removal with the IPAC method, tech. rep., IPAC Lab., Juillet 1998.

TWG2001 [62] P. García-Lario, A. Coulais, D. Kester, and E. Caux, Transients Working Group Final Report, ESA, Avril 2001.

Groezinger92ESA [63] U. Groezinger, S. Kirches, D. Lemke, J. Schubert, B. Schulz, and J. Wolf, Characterization of detectors and calibration of the ISOPHOT experiment, in ESA SP-356 Photon Detectors for Space Instrumentation, Décembre 1992, pp. 329–333.

HaegelBLH89 [64] N. M. Haegel, J. W. Beeman, P. N. Luke, and N. N. Haller, Transient photoconductivity in Ge:Be due to Be^{+} formation, Physical Review B, 39(6):3677–3682, Février 1989.

Haegel96 [65] N. M. Haegel, C. R. Brennan, and A. M. White, Transport in extrinsic photoconductors: a comprehensive model for transient response, Journal of Applied Physics, 80(3):1510–1514, Août 1996.

Haegel93 [66] N. M. Haegel, C. A. Latasa, and A. M. White, Transient response of infrared photoconductors : the roles of contacts and space charge, Applied Physics, A 56:15–21, 1993.

Haegel96proc [67] N. M. Haegel, C. Newton, J. C. Simoes, and A. M. White, Modeling of transient response in far infrared photoconductors, in Submillimeter and far infrared space instrumentation, 30th ES-LAB Symp, ed., no. SP-388, ESTEC, Décembre 1996, ESA, ESA, pp. 15–20.

Haegel99 [68] N. M. Haegel, J. C. Simoes, A. M. White, and J. W. Beeman, Transient behavior of infrared photoconductors: application of a numerical model, Applied Optics, 38(10):1910–1919, Avril 1999.

Haegel89 [69] N. M. Haegel and A. M. White, Modeling of near-contact field and carrier distributions in extrinsic germanium photoconductors, Infrared Physics, 29(5):915–923, 1989.

Hiromoto92 [70] N. Hiromoto, T. Itabe, H. Shibai, H. Matsuhara, T. Nakagawa, and H. Okuda, Threeelement stressed Ge:Ga photoconductor array for the infrared telescope in space, Applied Optics, 31:460–465, Février 1992.

IPAC1 [71] IPAC group, ISOCAM transient treatment at IPAC, tech. rep., IPAC Lab., 1997.

Kessler96 [72] M. F. Kessler et al., The Infrared Space Observatory (ISO) mission., Astron. Astrophys., 315:L27– L31, Novembre 1996.

Kester99 [73] D. Kester, The impact of memory effects correction on SWS data, tech. rep., SRON, Netherlands, 1999.

Lari97 [74] C. Lari, IRA model for ISOCAM LW transient correction, tech. rep., IRA, Juin 1997.

Lari2001 [75] C. Lari et al., A new method for ISOCAM data reduction - I. Application to the European Large Area ISO Survey Southern Field: method and results, MNRAS, 325:1173–1189, Août 2001.

PHThand99 [76] R. Laureijs, U. Klass, P. Richards, and B. Schulz, ISOPHOT Data Users Manual, V 4.1, ESA, Octobre 1999.

Leeks99 [77] S. J. Leeks, B. M. Swinyard, T. Lim, and P. E. Clegg, Aspects of LWS processing, The Universe as Seen by ISO. Eds. P. Cox & M. F. Kessler. ESA-SP 427., 427:81–84, Mars 1999.

Lemke96 [78] D. Lemke et al., ISOPHOT - capabilities and performance., Astron. Astrophys. Suppl. Ser., 315:L64–L70, Novembre 1996.

MAMD2000 [79] M.-A. Miville-Deschênes, F. Boulanger, A. Abergel, and J.-P. Bernard, Optimizing ISO-CAM data processing using spatial redundancy, Astron. Astrophys. Suppl. Ser., 146:519–530, Novembre 2000.

Moneti98 [80] A. Moneti, ed., ISO detector workshop, viewgraphs of the presentations, 14–16 Jan., vol. I–II, Villafranca del Castillo, 1998, ESA.

Mottier91 [81] P. Mottier, P. Agnese, and P. O. Lagage, Si:Ga focal plane arrays for satellite and ground-based telescopes, Proceedings of SPIE, 1494:419–426, Septembre 1991.

Mottier90 [82] P. Mottier, C. Lucas, M. Ravetto, and P. Agnese, Recent developments on ISOCAM long wavelength channel detector, Proceedings of SPIE, 1341:368–374, Novembre 1990.

Okumura98 [83] K. Okumura, ISOCAM PSF report, tech. rep., ESA, Vilspa, Juin 1998.

Okumura2000 [84] K. Okumura, The ISO point spread function and CAM profiles, in ISO beyound point sources: studies of extended infrared emission; ESA SP-455, Madrid, Spain, 14–17 Sept. 1999, Avril 2000.

Patra98b [85] A. I. Patrashin and G. A. Ivanov, Testing unit for measuring photoelectric characteristics of IR arrays, Proceedings of SPIE, 3379:555–560, Juillet 1998.

Patra97 [86] M. A. Patrashin, B. I. Fouks, U. Grozinger, D. Lemke, and J. Wolf, Residual conductivity of stressed Ge:Ga photoconductors after low-dose gamma irradiation, Journal of Applied Physics, 82(3):1450–1453, Août 1997.

1995SPIE.2475..476P [87] M. A. Patrashin, B. I. Fouks, D. Lemke, and J. Wolf, Radiation-induced effects in extrinsic far-infrared detectors, Proceedings of SPIE, 2475:476–486, Juin 1995.

Patra98a [88] M. A. Patrashin, M. Fujiwara, N. Hiromoto, and B. I. Fouks, Far-infrared Ge:Ga detector for low-background space-based astronomical observations, Proceedings of SPIE, 3354:232–239, Août 1998.

Perault94 [89] M. Pérault, F.-X. Désert, A. Abergel, F. Boulanger, C. Dupraz, A. Soufflot, C. Cesarsky, and L. Vigroux, Infrared Space Observatory camera calibration facility and preflight characterization, Optical Engineering, 33:762–770, Mars 1994.

NumRec92 [90] W. Press, S. Teukolsky, W. Vetterling, and B. Flannery, Numerical Recipes in FORTRAN, second edition, Cambridge University Press, 1992.

Schubert95PhD [91] J. Schubert, Die Eichung des Infrarot-Gitterspektrometers im Satelliten-Experiment ISOPHOT und die Korrektur der Transienten seiner Detektoren, PhD thesis, Max Planck-Institut für Astronomie, Heidelberg, 1995.

SchuFouks95 [92] J. Schubert, B. I. Fouks, D. Lemke, and J. Wolf, Transient response of ISOPHOT Si:Ga infrared photodetectors: experimental results and application of the theory of nonstationary processes, Proceedings of SPIE, 2553:461–469, Septembre 1995.

Schubert94 [93] J. Schubert, G. Roth, J. Wolf, D. Lemke, and B. I. Fouks, Correction and curing of inorbit-induced nonideal behavior of ISOPHOTS’s photodetectors, Proceedings of SPIE, 2268:283–294, Septembre 1994.

Sclar84 [94] N. Sclar, Properties of doped silicon and germanium infrared detectors, Prog. Quant. Electr., 9:149– 257, 1984.

See99 [95] J. Sée, Correction des effets mémoire de l’instrument ISOPHOT C-100, tech. rep., Stage de magistère, 1999.

SheinShi76 [96] M. K. Sheinkman and A. Y. Shik, Long-term relaxation and residual conductivity of semiconductors (review), Sov. Phys. Semicond., 10(2):128–143, Février 1976.

Siebenmorgen96 [97] R. Siebenmorgen et al., First results from the ISOCAM parallel mode, Astron. Astrophys., 315:L169–L172, Novembre 1996.

Starck99a [98] J.-L. Starck, A. Abergel, et al., ISOCAM data processing, Astron. Astrophys. Suppl. Ser., 134:135–148, Janvier 1999.

Starck99 [99] J.-L. Starck et al., Faint source detection in ISOCAM images, Astron. Astrophys. Suppl. Ser., 138:365–379, Août 1999.

SurisFouks78 [100] R. A. Suris and B. I. Fouks, Theory of transient processes in compensated semiconducteurs. I, Sov. Phys. Semicond., 12(12):1380–1384, Décembre 1978.

SurisFouks79 [101] R. A. Suris and B. I. Fouks, Theory of transient processes in compensated semiconducteurs. II, Sov. Phys. Semicond., 13(1):79–86, Janvier 1979.

SurisFouks80N1 [102] R. A. Suris and B. I. Fouks, Noise in finite compensated semiconductors. I. Generationrecombinaison noise, Sov. Phys. Semicond., 14(6):641–646, Juin 1980.

SurisFouks80N2 [103] R. A. Suris and B. I. Fouks, Noise in finite compensated semiconductors. II. johnson noise and contact noise, Sov. Phys. Semicond., 14(6):647–652, Juin 1980.

SurisFouks80 [104] R. A. Suris and B. I. Fouks, Theory of nonlinear processes in compensated semiconducteurs, Sov. Phys. Semicond., 14(8):896–901, Août 1980.

Swinyard98tech [105] B. Swinyard, P. Clegg, S. Leeks, M. Griffin, T. Lim, and M. Burgdorf, Space operation and performance of doped germanium photoconducting detectors in the far infrared: experience from the ISO LWS, tech. rep., ESA, 1998.

Swinyard1998SPIE [106] B. M. Swinyard, M. J. Burgdorf, P. E. Clegg, G. R. Davis, M. J. Griffin, C. Gry, S. J. Leeks, T. L. Lim, S. Pezzuto, and E. Tommasi, In-orbit performance of the ISO Long-Wavelength Spectrometer, Proceedings of SPIE, 3354:888–899, Août 1998.

Swinyard96 [107] B. M. Swinyard et al., Calibration and performance of the ISO Long-Wavelength Spectrometer., Astron. Astrophys. Suppl. Ser., 315:L43–L48, Novembre 1996.

TeitLW86 [108] S. W. Teitsworth, B. A. Lynn, and R. M. Westervelt, Chaotic dynamics in a simple model of extrinsic photoconductors, Physical review letters, T 14:71–75, 1986.

TeitWest84 [109] S. W. Teitsworth and R. M. Westervelt, Chaos and broadband noise in extrinsic photoconductors, Physical review letters, 53(27):2587–2590, Décembre 1984.

Vigroux91 [110] L. Vigroux, ISOCAM : User’s manual, tech. rep., C.E.A. CEN Saclay, DSM, S.Ap., Janvier 1991.

Vigroux93 [111] L. G. Vigroux, C. J. Cesarsky, O. Boulade, Y. Rio, M. Perault, A. Abergel, F.-X. Désert, D. Rouan, and F. Lacombe, ISOCAM–the Infrared Space Observatory camera: results of testing and calibrations, Proceedings of SPIE, 1946:281–292, Octobre 1993.

VinoFouks88 [112] L. A. Vinokurov and B. I. Fouks, Spreading of the current in a photosensitive medium exhibiting an extrinsic photoconductivity, Sov. Phys. Semicond., 22(11):1258–1262, Novembre 1988.

VinoFouks91 [113] L. A. Vinokurov and B. I. Fouks, Nonlinear photoresponse of extrinsic photoconductors, Sov. Phys. Semicond., 25(11):1207–1211, Novembre 1991.

VinoZKF92 [114] L. A. Vinokurov, N. B. Zaletaev, V. F. Kocherov, and B. I. Fouks, Spreading of time dependent photocurrrent in Si:Ga, Sov. Phys. Semicond., 26(4):433–435, Avril 1992.

Wensink92ESA [115] J. W. Wensink et al., Characteristics of the ISO Short-Wavelength Spectrometer detectors, in ESA SP-356 Photon Detectors for Space Instrumentation, Décembre 1992, pp. 339–344.

WestTeit85 [116] R. M. Westervelt and S. W. Teitsworth, Nonlinear transient response of extrinsic Ge farinfrared photodetectors, Journal of Applied Physics, 57(2):5457–5469, Juin 1985.

White85 [117] A. M. White, The characteristics of monority-carrier exclusion in narrow direct gap semoconductors, Infrared Physics, 26(6):729–741, 1985.

Wolf94 [118] J. Wolf, Low-background far-infrared detectors and arrays, Optical Engineering, 33(5):1492–1500, Mai 1994.